An 18kg object is pulled by a force of 210N, 40° with the horizontal. The coefficient of kinetic friction of the object is given as 0.45. What is the acceleration of the object?


1
Expert's answer
2021-12-14T15:08:41-0500

"210\cdot\cos40°-\mu N=ma"


"N+210\cdot\sin40^°-mg=0\to N=mg-210\cdot\sin40°="


"=18\cdot9.8-210\cdot\sin40°=41.4\ (N)"


"a=(210\cdot\cos40°-0.45\cdot41.4)/18=7.9\ (m/s^2)"